# Local Order Average-Atom Interatomic Potentials

### v2

@@ III.2.3 Phase Transformations in NiAl
Figure 20: All images are TOP view, looking along $z$. (a) Transformed structure of the LOAA simulation at the end of the cooling phase (the initial high temperature phase was the same as Fig. 17(b)). In (b), we see that the structure is still fct martensite at about 620K during heating, just before the first rise in the heating curve. (c) The structure at about 700K, just after the first rise, which we see by the direction of the shear is still fct, but a different variant. (d) The final structure at 900K, which has transformed back to B2.
Using the AA formulation produces essentially the same behavior as the random TS sample, i.e. there is a spontaneous transformation to the fcc phase because the B2 phase is unstable for the random alloy. This is shown in Fig. 18. On the other hand, the LOAA formulation shows the same B2 stability at high temperatures and the phase transformation as the ordered TS sample, albeit with different transformation temperatures. This can be seen in Figs. 19 and 20. The hysteresis loop for LOAA (Fig. 19) shows an unusual double rise on heating. The first rise at around 650 K is actually not the phase transition, but rather a jump between two variants of the martensite before the final true transformation to the austenite at around 755 K. This is shown in the snapshots of Fig. 20, where (a) is the transformed martensite after cooling, (b) is the same martensite but at a temperature of around 620K, just before the first abrupt rise in the heating curve. In (c), we see that the crystal is still in the martensitic phase at about 640 K (after the rise), although sheared in the other direction. Finally, (d) shows the crystal after it has fully transformed back to B2 austenite, made clear by the orthogonal simulation box and the blue coloring (the noisy mix of CNA values in (b)–(d) is due to atomic vibrations at high temperature).
@@ IV Summary
An extension to the AA approach accounting for LO (short-range order) is derived based on partial RDF information obtained on a reference configuration. The result is an effective IP defined for an atomistic system without species, which in the limit of no structural relaxation provides exact results for an ensemble over all possible species arrangements on the reference configuration. Once the effective IP is constructed, it has a similar cost to the original “true species” IP upon which it is based, but enables simulations with far smaller supercells and without the need to run multiple times for good statistics for significant computational savings.
LOAA is validated on several example problems. First, a 2D binary hexagonal crystal modeled via LJ interactions. A phase diagram is constructed showing the dependence of the nature and intensity of LO as a function of LJ parameters. It is shown that LOAA accurately reproduces the energy and elastic constants of the true species IP for LO patterns, in contrast to AA that is only correct for the random alloy case where no LO exists. Second, LOAA is shown to accurately reproduce the lattice constant, cohesive energy, and elastic properties of an ordered Fe(1-x)/2Ni(1-x)/2Crx alloy across for $x$ in the range 0 to 0.5 modeled via an EAM potential. Finally, third, a temperature driven martensitic phase transformation under axial compression in an off-stoichiometry Ni0.67Al0.33 alloy modeled via EAM is studied. It is shown that LO is necessary to stabilize the high-temperature B2 phase. Upon cooling the crystal exhibits hysteresis as it transforms to a low temperature fct phase, and then back to B2 upon heating. The LOAA effective IP is able to capture this effect, albeit at somewhat different transformation temperatures, whereas the AA effective destabilizes into a twinned fcc-like structure.
The method can be readily extended to other IP functional forms, and would be of interest to apply to complex materials, like HEAs that require very large supercells to capture the off-stoichiometric structure, and where LO is known to play an important role. On the methodological side, the approach can be extended to account for local relaxation effects by considering the effect of such relaxation on the partial RDF form. Preliminary work along these lines appears promising.
Code and example input suitable for running LOAA in LAMMPS is provided in the SM.
@@ Acknowledgements.
CZ and ET acknowledge support by NASA Space Technology Graduate Research Opportunity (NSTGRO) Award 21-NSTGRO22-0159. RM acknowledges the support of the Natural Sciences and Engineering Research Council of Canada (NSERC). The authors thank Ryan S. Elliott for helpful discussions, and Lauren Abbott and the computational materials group at NASA Ames Resarch Center for their valuable assistance.
@@ Appendix B Relative bonding energies in 2D Hexagonal Binary Crystals
Appendix C Theoretical Analysis of the Error in Energy Predictions of the AA Method
@@
Appendix D Elastic Constant Equations for AA and LOAA
@@
The motivation for this substitution will be more evident in the LOAA formulation below. Following the same change of variables procedure in Appendix C, this can be re-written as

### v4

@@ III.2.2 Phase Transformations in NiAl
Figure 13: Snapshots of the Ni0.67Al0.33 system computed using LOAA in TOP view looking along $z$. (a) Transformed structure of the LOAA simulation at the end of the cooling phase (the initial high temperature phase was the same as Fig. 10(b)). In (b), we see that the structure is still fct martensite at about 620K during heating, just before the first rise in the heating curve. (c) The structure at about 700K, just after the first rise, which we see by the direction of the shear is still fct, but a different variant. (d) The final structure at 900K, which has transformed back to B2.
Using the AA formulation produces essentially the same behavior as the random TS sample, i.e. there is a spontaneous transformation to the fcc phase because the B2 phase is unstable for the random alloy. This is shown in Fig. 11. On the other hand, the LOAA formulation shows the same B2 stability at high temperatures and the same phase transformation as the ordered TS sample, albeit with different transformation temperatures. This can be seen in Figs. 12 and 13. The hysteresis loop for LOAA (Fig. 12) shows an unusual double rise on heating. The first rise at around 650 K is actually not the phase transition, but rather a jump between two variants of the martensite before the final true transformation to the austenite at around 755 K. This is shown in the snapshots of Fig. 13, where (a) is the transformed martensite after cooling, (b) is the same martensite but at a temperature of around 620K, just before the first abrupt rise in the heating curve. In (c), we see that the crystal is still in the martensitic phase at about 640 K (after the rise), although sheared in the other direction. Finally, (d) shows the crystal after it has fully transformed back to B2 austenite, made clear by the orthogonal simulation box and the blue coloring (the noisy mix of CNA values in (b)–(d) is due to atomic vibrations at high temperature).
@@@ GOT TO HERE @@@
@@ IV Summary
LOAA, an extension to the AA approach accounting for LO (short-range order), is derived based on partial RDF information obtained on a reference configuration. The result is an effective IP defined for an atomistic system without species that in the limit of no structural relaxation provides exact results for an ensemble over all possible species arrangements on the reference configuration. Once the effective IP is constructed, it has a similar cost to the original “true species” IP upon which it is based, but enables simulations with far smaller supercells and without the need to run multiple times for good statistics, leading to significant computational savings. The LOAA method is attractive for simulations of complex materials, such as HEAs that require very large supercells to capture the off-stoichiometric structure, and where LO is known to play an important role.
The LOAA potential form is derived here for pair and EAM-style potentials, and can be extended to other IP functional forms in a similar fashion. An interesting observation is that LO effects only enter EAM-style potentials through the pairwise term. This is due to the assumption of charge spherical symmetry in EAM that implies that the charge density at an atomic site depends separately on the identities of the contributing atoms (and not on pairs of atom identities).
LOAA is validated on several example problems. First, a 2D binary hexagonal crystal modeled via LJ interactions. A phase diagram is constructed showing the dependence of the nature and intensity of LO on the LJ parameters. It is shown that LOAA accurately reproduces the energy and elastic constants of the true species IP across the phase diagram, in contrast to AA that is only correct for the random alloy case where no LO exists. Second, LOAA is shown to accurately reproduce the lattice constant, cohesive energy, and elastic properties of an ordered Fe(1-x)/2Ni(1-x)/2Crx alloy for $x$ in the range 0 to 0.5 modeled via an EAM potential. Third, a temperature driven martensitic phase transformation under axial compression in an off-stoichiometry Ni0.67Al0.33 alloy modeled via EAM is studied. It is shown that LO is necessary to stabilize the high-temperature B2 phase. Upon cooling the crystal exhibits hysteresis as it transforms to a low temperature fct phase, and then back to B2 upon heating. The LOAA effective IP is able to capture this effect, albeit at somewhat different transformation temperatures, whereas in AA simulations the alloy destabilizes into a twinned fcc-like structure.
Code and example input suitable for running LOAA in LAMMPS is provided in the SM. Additional validation of results is also provided in the SM.
@@ Acknowledgements.
CZ and ET acknowledge support by NASA Space Technology Graduate Research Opportunity (NSTGRO) Award 21-NSTGRO22-0159. RM acknowledges the support of the Natural Sciences and Engineering Research Council of Canada (NSERC). The authors thank Ryan S. Elliott for helpful discussions, and Lauren Abbott and the computational materials group at NASA Ames Research Center for their valuable assistance.
@@ Appendix B Relative bonding energies in 2D Hexagonal Binary Crystals
Appendix C Understanding the ‘no-LO’ Line of the Phase Diagram
The line in the phase diagram where no local-ordering occurs (i.e. the ‘no-LO line’) is defined by the equation:
Appendix D Theoretical Analysis of the Error in Energy Predictions of the AA Method
@@
Appendix E Elastic Constant Equations for AA and LOAA
@@
The motivation for this substitution will be more evident in the LOAA formulation below. Following the same change of variables procedure in Appendix D, this can be re-written as
@@
We can derive expressions for the slopes of the elastic constants when plotted against $t$ be separating out the terms multiplying $t$,